OpenStudy (josedavid):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 20

OpenStudy (anonymous):

equation of a circle with center (a,b) and radius r is \[(x-a)^2+(y-b)^2 = r^2\] can you try rearranging the equation ?

OpenStudy (josedavid):

what do u mean

OpenStudy (anonymous):

Try rearranging the given equation to the equation of a circle as mentioned above in my reply. \[x^2+2x+y^2+4y = 20\] => \[x^2-2x(-1))+(-1)^2+y^2+2*y*(2)+(2)^2 = 25\] => \[(x-(-1))^2 + (y-(-2))^2 = (5)^2\] radius = 5 and center = (-1,-2)

OpenStudy (josedavid):

what is the answer what do i write